OpenStudy (ashsmashum1290):

What is the approximate circumference of the circle? A.1.57 m B.157 m C.314 m D.3.14 m

OpenStudy (ashsmashum1290):



TheSmartOne (thesmartone):

\(\sf Circumference = 2 \pi r\) r in your diagram is .5. And \(\sf \pi = 3.14\) Plug it in and solve.

OpenStudy (ashsmashum1290):

1.57

OpenStudy (ashsmashum1290):

what do you guys think?

OpenStudy (nathanjhw):

2pi r = 2(3.14) (0.5)

OpenStudy (nathanjhw):

Plug that into a calculator.

OpenStudy (ashsmashum1290):

1.57?

OpenStudy (nathanjhw):

No. Did you plug it into a calculator.

OpenStudy (ashsmashum1290):

yeah

OpenStudy (nathanjhw):

To make it easier, perhaps you won't need a calculator, 0.5 is also equal to dividing by 2. Therefore you can write the equation as

OpenStudy (nathanjhw):

\[2\pi/2 \]

OpenStudy (nathanjhw):

The 2's cross out and you are left with?

OpenStudy (ashsmashum1290):

3.14

OpenStudy (nathanjhw):

Yes.

TheSmartOne (thesmartone):

Good job at explaining it @NathanJHW :)

OpenStudy (ashsmashum1290):

so its D

OpenStudy (nathanjhw):

Exactly.

OpenStudy (ashsmashum1290):

Thanks @NathanJHW

OpenStudy (nathanjhw):

You're welcome.